# Bivariate phase-type distributions for experience rating in disability insurance

### v2

@@ 2.4.2. Numerical considerations
When working with a large volume of data points, quantities such as $y_{\bullet j}=\sum_{i=1}^{d}y_{ij}$ can take considerable large values, making the computation of some of the previous expressions unfeasible in the form given. However, this can be easily solved by working with alternative and equivalent expressions. More specifically, for the joint density, it is convenient to work with
@@
3. Experience rating for disability insurance
We seek to apply the methodology surveyed and developed in the previous section to price group disability insurance contracts subject to heterogeneity. Subsection 3.1 contains a description of the insurance coverage of interest. In Subsection 3.2, we consider shrinkage estimation of disability and recovery rates subject to, among other things, right-censoring and following closely Furrer, (2019). In particular, we provide a link to the mixed Poisson models of Section 2.
3.1. Disability coverage
$U(t)=t-\sup\{s\in[0,t]:Z_{s}\neq Z_{t}\}.$
Finally, $\varepsilon\in[0,\infty)$ is a waiting period, $\tau\in(0,\infty]$ is the coverage period, $t\mapsto b_{ai}(t)$ are deterministic benefits upon disability, and $t\mapsto b_{i}(t)$ is a deterministic benefit rate during disability.
Example 3.1.
In the following, we suppose that the payment functions $b_{ai}$ and $b_{i}$ are suitably regular and that there exists a maximal contract time $T\in(0,\infty)$ such that $b_{ai}(t)=b_{i}(t)=0$ for $t>T$. This is always the case in practice (with $T$ related to, for instance, the legal retirement age). Furthermore, we suppose that the initial state of $Z$ is $a$, and we naturally take $d$ to be absorbing.
We further assume that $Z$ is Markovian and admits transition rates $t\mapsto\mu_{jk}(t)$, $j,k\in\mathcal{S}$, $j\neq k$, such that
$\mu_{jk}(t)=\lim_{h\downarrow 0}\frac{\mathbb{P}(Z_{t+h}=k\,|\,Z_{t}=j)}{h}$
In other words, we have
The model is illustrated in Figure 3.1.
[…]
where $t\mapsto r(t)$ is a suitably regular interest rate of choice. We are, of course, implicitly requiring that $s\leq t$. The following result is a consequence of, for instance, Corollary 7.2 in Adékambi and Christiansen, (2017).
Proposition 3.2.
It holds almost everywhere that
Section 8 in Adékambi and Christiansen, (2017) provides a numerical scheme for the prospective reserves based on (3.2) in conjunction with the obvious boundary conditions $V_{a}(T)=0$ and $V_{i}(T,s)=0$.
Remark 3.1.
This leads to a particularly straightforward numerical scheme.
3.2. Shrinkage estimation
Experience rating for group disability insurance concerns shrinkage estimation of group effects. To this end, Furrer, (2019) suggests estimation of the transition rates using empirical Bayes methods. Disregarding mortality, the ‘simple shrinkage’ estimation of Furrer, (2019) is based on the following structural assumptions:
$O_{ai}^{k}\,|\,(\Theta_{ai},\Theta_{ia})$ $\sim\text{Poisson}\big{(}\Theta_{ai}E_{a}^{k}\mu_{ai}(t_{k})\big{)},$
$O_{ia}^{k}\,|\,(\Theta_{ai},\Theta_{ia})$ $\sim\text{Poisson}\big{(}\Theta_{ia}E_{i}^{k}\mu_{ia}(t_{k})\big{)}.$
This provides a direct link to Section 2.
The next section contains a thorough numerical study devoted to experience rating for disability insurance, building on this and the previous section. We adopt the empirical Bayes and ‘plug-in’ approach of Furrer, (2019), but study other types of mixing than that of (3.5), namely the hierarchical and phase-type mixings of Section 2. By allowing for dependence between latent group effects, estimation of the recovery rate may borrow strength from the estimation of the disability rate – and vice versa. Everything else being equal, this should lead to improvements in predictive performance.
@@ 4. Numerical study
We now use simulated data to provide a detailed example of the practical implementation of the proposed methodologies for experience rating for disability insurance. We focus on the disability model depicted in Figure 3.1 with group effects. In particular, we delve into the shrinkage estimation of the disability and recovery rates. This hands-on approach will not only demonstrate the feasibility of the techniques proposed in the previous sections, but also provide a concrete understanding of their performance. All simulations and numerical computations were performed using the R programming language.
The numerical study is divided into three parts. In Subsection 4.1, we describe the methodology and approach taken to simulate an insurance portfolio with different groups, providing insights into the underlying assumptions and scenarios. Subsection 4.2 focuses on the outcomes of the estimation of the models under study, presenting a detailed examination of the results and drawing comparisons between them. Finally, in Subsection 4.3, we explore the predictive performance of the models by examining the results in various iterations of the simulation study and assessing their effectiveness for pricing.
@@ 4.1. Simulation
We begin our numerical study by simulating an insurance portfolio comprised of $100$ distinct groups and a total of $50{,}000$ insured, each having a coverage period of $3$ years. Recognizing that real-world portfolios often comprise groups of varying sizes, we have allocated the insured to groups of differing sizes, confer with Figure 4.1. The group sizes follow, approximately, a Gamma distribution with shape parameter $0.05$ and rate parameter $0.0001$.
@@
Next, we focus on the initial age composition within each group. To model the initial ages of the different groups, we randomly generate entry ages between $20$ and $67$ years following a mixture distribution. The choice of a mixture distribution is to reflect different age profiles in the portfolio, while the age range is representative of an actual insurance portfolio. More specifically, for the mixture distribution, we consider the following three distinct mixture components: A Weibull distribution with shape parameter $3$ and scale parameter $15$, representing younger groups, a Normal distribution with mean $47$ and standard deviation $7$, symbolizing middle-aged groups, and a Gompertz distribution with shape parameter $0.1$ and rate parameter $0.002$, characterizing older groups. The corresponding mixture weights are $0.25$, $0.35$, and $0.4$, respectively. Figure 4.2 depicts the described mixture distribution along with a visual representation of the simulated initial ages.
@@
Figure 4.3 shows the transition rates just defined.
@@
It is worth mentioning that we keep the above-described initial group sizes, age distributions, and baseline transition rates fixed throughout the study.
@@
One of the primary goals of this study is to investigate how the proposed models perform with respect to different dependence structures of the mixing vector $\bm{\bm{\Theta}}$. Thus, we consider three possible scenarios: independence, positive correlation, and negative correlation. This setting, along with the two distinct marginal specifications, enables us to evaluate the robustness and flexibility of the reviewed models in capturing complex joint behaviors, reflecting various possible real-world situations for disability insurance portfolios. More specifically, to introduce dependency within the random vector $\bm{\bm{\Theta}}$, we employ a Clayton copula with three different values of Kendall’s tau, $\rho_{\tau}$: $-0.5$, 00, and $0.5$, thus encompassing all three dependency cases.
@@ 4.2. Estimation
Having simulated a dataset for Case A, Case B, and Case C, respectively, we now proceed to perform shrinkage estimation of the disability and recovery rates using the independent Gamma, which we also call ‘simple’ following Furrer, (2019), the hierarchical Gamma, and the bivariate phase-type model. It is essential to note that we set $p_{1}=p_{2}=6$ for the bivariate phase-type mixing to allow for enough flexibility to capture the different dependence structures considered. Across all models, we assume the following parametric form of the baseline transition rates:
@@
Recall that shrinkage estimation consists of first performing maximum likelihood estimation of the corresponding mixed Poisson regression models via the EM algorithms presented in Section 2, thus obtaining different estimators of the regression parameters, ${\boldsymbol{\beta}}^{(ai)}$ and ${\boldsymbol{\beta}}^{(ia)}$, and the parameters of the specifications of $\bm{\bm{\Theta}}$. Then, the shrinkage estimates of $\bm{\bm{\Theta}}$ are obtained by computing the Bayes estimators under quadratic loss utilizing the previously computed maximum likelihood estimates. Finally, if we denote by $\hat{\bm{\bm{\Theta}}}^{(g)}=(\hat{\Theta}_{ai}^{(g)},\hat{\Theta}_{ia}^{(g)})$ the resulting shrinkage estimates for each group, $g=1,\dots,100$, and by $\hat{\boldsymbol{\beta}}^{(ai)}$ and $\hat{\boldsymbol{\beta}}^{(ia)}$ the estimates of the regression parameters, the ‘plug-in’ approach uses as estimates of the disability and recovery rates exactly
@@
For a visual comparison of the fitted models, we would like to begin by providing plots of the original (simulated) group effects against the estimated ones. Nevertheless, given the presence of an intercept in the regression and also due to the well-known identifiability issue of phase-type distributions (confer with Yslas,, 2021), it becomes more suitable to graph the original group effects against the following scaled estimates:
@@
where ${\boldsymbol{\beta}}_{\text{true}}^{(ai)}=(-4.5,-0.018,0.00064)$, ${\boldsymbol{\beta}}_{\text{true}}^{(ai)}=(0.3,-0.049)$, and $E_{a}^{k,(g)}$ and $E_{i}^{k,(g)}$ are the relevant exposures across groups $g=1,\dots,100$. This procedure allows for a more meaningful comparison as it mitigates the aforementioned effects. The resulting plots for the group effects of the disability and recovery rates for Case A can be found in Figure 4.5, while Figure 4.6 contains the corresponding illustrations for Case B. We observe that the models specifying a random effect tend to outperform the fixed effect one. This performance difference is particularly noticeable in the cases of groups without observed disabilities, leading to an estimation of a group effect equal to zero in the fixed effect case, while this is not the case for the random effects models. However, from the plots alone, it is not evident which model for $\bm{\bm{\Theta}}$ performs best in each case. Therefore, we now turn our attention to the analysis of the estimates of the transition rates. We focus on the groups $16$, $23$, and $68$.
@@
Figure 4.5. Estimated group effects against true (simulated) values for Case A.
Figure 4.6. Estimated group effects against true (simulated) values for Case B.
@@
For Case A, the estimates of the disability rates can be found in Figure 4.7, while the corresponding recovery rates are illustrated in Figure 4.8. For the disability rates (Figure 4.7), while the estimates of group $16$ are similar across all models considered, for the other two groups ($23$ and $68$), the three shrinkage fits outperform with similar results the standard and fixed effect fits. The superiority of the shrinkage estimates becomes more evident in the recovery rates (Figure 4.8), where they result in closer approximations to the true rates. This is particularly clear for group $23$, where the fixed effect model estimates a group effect of zero. We also observe the downside of the standard model, which yields identical estimates for all groups. Interestingly, in the case of the largest of the three groups, group $68$, the hierarchical model is outperformed by the simple and phase-type models. This may be attributed to the hierarchical model’s limitation in capturing independence between the random effects.
@@
Moving on to Case B, estimates of the disability rates are depicted in Figure 4.9, and a similar conclusion can be drawn to that of Case A, with the shrinkage estimates showing a clear advantage over the standard and fixed effect fits. However, the primary distinction among the shrinkage estimates is observable in the recovery rates, presented in Figure 4.10. More specifically, for groups $16$ and $68$, the models capable of capturing positive correlation – namely, the hierarchical and phase-type models – provide closer approximations to the true disability rates. This leads us to conclude that by allowing dependency between latent group effects, the estimates of recovery and disability rates can mutually benefit.
@@
Figure 4.7. Estimates of disability rates for the three selected groups in Case A.
@@
Figure 4.8. Estimates of recovery rates for the three selected groups in Case A.
Figure 4.9. Estimates of disability rates for the three selected groups in Case B.
Figure 4.10. Estimates of recovery rates for the three selected groups in Case B.
@@
We conclude this part of the analysis by presenting the loglikelihoods of the fitted mixed Poisson regression models. These loglikelihoods serve as statistical means to further assess the quality of the fits and can be found in Table 4.2. The observations from the table are consistent with what one could expect. In Case A, the simple model outperforms others, with the phase-type model following closely. This result aligns with the nature of the models and the data considered for this case. For Case B, the hierarchical model yields the highest loglikelihood, followed closely by the phase-type model. This outcome is somewhat anticipated since these two models are capable of capturing the positive correlation present in the mixing vector, aligning well with the underlying structure. Finally, for Case C, the phase-type model takes the lead, surpassing the other two models. This is hardly surprising, as it is the only model that can capture a negative correlation. Overall, the phase-type model proves to be very competitive across all cases. Its flexibility to capture different dependence structures highlight its potential as a robust tool for experience rating modeling.
@@
Simple $-3924.66$ $-3816.12$ $-3639.01$
Hierarchical $-3930.41$ $-3804.25$ $-3657.02$
Phase-type $-3925.29$ $-3805.56$ $-3627.30$

### v3

@@ 3.4.2. Numerical considerations
When working with a large volume of data points, quantities such as $y_{\bullet j}=\sum_{i=1}^{d}y_{ij}$ can take considerably large values, and the direct evaluation of the previous expressions, which involve factorials, powers, and exponentials, may lead to combinations of extremely large and small numbers, making the computation of some of the previous expressions unfeasible in the form given. However, this can be easily solved by working with alternative and equivalent expressions, where such terms combine to result in numerically stable quantities. More specifically, for the joint density, it is convenient to work with
@@ 4. Numerical study
We now use simulated data to provide a detailed example of the practical implementation of the proposed methodologies for experience rating for disability insurance. We focus on the disability model depicted in Figure 2.1 with group effects. In particular, we delve into the shrinkage estimation of the disability and recovery rates. This hands-on approach will not only demonstrate the feasibility of the techniques proposed in the previous sections, but also provide a concrete understanding of their performance. All simulations and numerical computations were performed using the R programming language, with the corresponding code available through the GitHub repository https://github.com/jorgeyslas/bivph-experience-rating.git.
The numerical study is divided into three parts. In Subsection 4.1, we describe the methodology and approach taken to simulate an insurance portfolio with different groups, providing insights into the underlying assumptions and scenarios. Subsection 4.2 focuses on the outcomes of the estimation of the models under study, presenting a detailed examination of the results and drawing comparisons between them. Finally, in Subsection 4.3, we explore the predictive performance of the models by examining the results in various iterations of the simulation study and assessing their effectiveness for pricing. The insights from the last subsection are particularly important from a pratical point of view.
@@ 4.1. Simulation
We begin our numerical study by simulating an insurance portfolio comprised of $100$ distinct groups and a total of 50,00050{,}000 insured, each having a coverage period of $3$ years, but also a retirement age (maximal contract time) of $67$. Recognizing that real-world portfolios often comprise groups of varying sizes, we have – partly by hand – allocated the insured to groups of differing sizes, confer with Figure 4.1. The group sizes seem to follow, approximately, a Gamma distribution with shape parameter of about $0.05$ and rate parameter of about $0.0001$.
@@
Next, we focus on the initial age composition within each group. To model the initial ages of the different groups, we randomly generate entry ages between $20$ and $67$ years following a mixture distribution. The choice of a mixture distribution is to reflect different age profiles in the portfolio, while the age range is representative of an actual insurance portfolio. More specifically, for the mixture distribution, we consider the following three distinct mixture components: A Weibull distribution with shape parameter $3$ and scale parameter $15$, representing younger groups, a Normal distribution with mean $47$ and standard deviation $7$, symbolizing middle-aged groups, and a Gompertz distribution with shape parameter $0.1$ and rate parameter $0.002$, characterizing older groups. The corresponding mixture weights are $0.25$, $0.35$, and $0.4$, respectively. Ages are drawn from the mixture distribution conditional on the range from $20$ to $67$ years (via rejection and resampling). Figure 4.2 depicts the described mixture distribution along with a visual representation of the simulated initial ages.
@@
Figure 4.3 shows the transition rates just defined. It is worth mentioning that we keep the above-described initial group sizes, age distributions, and baseline transition rates fixed throughout the study.
@@
One of the primary goals of this study is to investigate how the proposed models perform with respect to different dependence structures of the mixing vector $\bm{\bm{\Theta}}$. Thus, we consider three possible scenarios: independence, positive correlation, and negative correlation. This setting, along with the two distinct marginal specifications, enables us to evaluate the robustness and flexibility of the reviewed models in capturing complex joint behaviors, reflecting various possible real-world situations for disability insurance portfolios. More specifically, to introduce dependency within the random vector $\bm{\bm{\Theta}}$, we employ a Clayton copula with three different values of Kendall’s tau, $\rho_{\tau}$: 0.5-0.5, 0, and $0.5$, thus encompassing all three dependency cases.
@@
Note that only Case A coincides with one of the considered models (independent Gamma). In Cases B and C, none of the models under comparison matches the true specification, in particular because the dependence structure is generated using a Clayton copula. This choice ensures that the models’ performance is assessed outside their own families as to not introduce unwanted bias in favor of one of the models. However, it is essential to point out that we purposely chose the same distribution for both margins, as the hierarchical model is restricted to having equal marginal variances. As shown at the end of the study, this restriction represents a significant limitation for the predictive performance of the hierarchical specification.
@@
Throughout, we keep the simulated group effects fixed – and the various models are compared in their ability to assess these simulated effects. Consequently, to understand the performance of the different models across scenarios, one needs to have in mind the empirical, rather than the underlying, distribution of these effects. Therefore, we report in Table 4.2 empirical estimates of the mean, standard deviation, and Kendall’s $\rho_{\tau}$ across cases.
Empirical $\Theta_{ai}$ (disability rate) $\Theta_{ia}$ (recovery rate)
Case A Case B Case C Case A Case B Case C
Mean $0.994437$ $0.853114$ $\phantom{-}0.988569$ $1.003830$ $0.916206$ $\phantom{-}0.955082$
Std. deviation $0.612241$ $0.584568$ $\phantom{-}0.552639$ $0.594999$ $0.512831$ $\phantom{-}0.476886$
Kendall’s $\rho_{\tau}$ $0.061818$ $0.565657$ 0.452525-0.452525 $0.061818$ $0.565657$ 0.452525-0.452525
Table 4.2. Empirical mean, standard deviation, and Kendall’s $\tau$ for the simulated group effects.
@@ 4.2. Estimation
Having simulated a dataset for Case A, Case B, and Case C, respectively, we now proceed to perform shrinkage estimation of the disability and recovery rates using the independent Gamma, which we also call ‘simple’ following Furrer, (2019), the hierarchical Gamma, and the bivariate phase-type model. It is essential to note that we set $p_{1}=p_{2}=3$ for the bivariate phase-type mixing in Case A, while for Cases B and C we use $p_{1}=p_{2}=6$. These dimensions were chosen to provide sufficient flexibility to capture the different dependence structures considered – particularly in Cases B and C – while keeping the estimation computationally feasible and ensuring stability in the subsequent predictive performance study based on multiple portfolio realizations. For practical applications, however, one may follow the standard model selection strategy used in phase-type fitting, beginning with low-dimensional specifications and gradually increasing complexity, with improvements assessed through, for instance, loglikelihood comparisons, which can be additionally supported with visual diagnostic tools (see Yslas,, 2025, for a detailed discussion).
Across all models, we assume the following parametric form of the baseline transition rates:
@@
Recall that shrinkage estimation consists of first performing maximum likelihood estimation of the corresponding mixed Poisson regression models via the EM algorithms presented in Section 3, thus obtaining different estimators of the regression parameters, ${\boldsymbol{\beta}}^{(ai)}$ and ${\boldsymbol{\beta}}^{(ia)}$, and the parameters of the specifications of $\bm{\bm{\Theta}}$. As it is customary in any form of EM algorithm, the number of steps is determined such that the absolute change in the parameter estimates or loglikelihoods between successive iterations becomes negligible. Then, the shrinkage estimates of $\bm{\bm{\Theta}}$ are obtained by computing the Bayes estimators under quadratic loss utilizing the previously computed maximum likelihood estimates. Finally, if we denote by $\hat{\bm{\bm{\Theta}}}^{(g)}=(\hat{\Theta}_{ai}^{(g)},\hat{\Theta}_{ia}^{(g)})$ the resulting shrinkage estimates for each group, $g=1,\dots,100$, and by $\hat{\boldsymbol{\beta}}^{(ai)}$ and $\hat{\boldsymbol{\beta}}^{(ia)}$ the estimates of the regression parameters, the ‘plug-in’ approach uses as estimates of the disability and recovery rates exactly
@@
For a visual comparison of the fitted models, we would like to begin by providing plots of the original (simulated) group effects against the estimated ones. Nevertheless, given the presence of an intercept in the regression and also due to the well-known identifiability issue of phase-type distributions (confer with Yslas,, 2025), it becomes more suitable to graph the original group effects against the following scaled estimates:
@@
where ${\boldsymbol{\beta}}_{\text{true}}^{(ai)}=(-4.5,-0.018,0.00064)$, ${\boldsymbol{\beta}}_{\text{true}}^{(ai)}=(0.3,-0.049)$, and $E_{a}^{k,(g)}$ and $E_{i}^{k,(g)}$ are the relevant exposures across groups $g=1,\dots,100$. This procedure allows for a more meaningful comparison as it mitigates the aforementioned effects. The resulting plots for the group effects of the disability and recovery rates across Cases A, B, and C are collected in Figure 4.5, where we have omitted the hierarchical model in Case C, since it only allows for positive correlation and therefore cannot be meaningfully compared in a setting with negative dependence. For this reason, the hierarchical specification is consistently excluded from the Case C analysis throughout the study. We observe that the models specifying a random effect tend to outperform the fixed effect one. This performance difference is particularly noticeable in the cases of groups without observed disabilities, leading to an estimation of a group effect equal to zero in the fixed effect case, while this is not the case for the random effects models. However, from the plots alone, it is not evident which model for $\bm{\bm{\Theta}}$ performs best in each case. Therefore, we now turn our attention to the analysis of the estimates of the transition rates. We focus on the groups $16$, $23$, and $68$.
@@
Figure 4.5. Estimated group effects against true (simulated) values.
@@
For Case A, the estimates of the disability rates can be found in Figure 4.9, while the corresponding recovery rates are illustrated in Figure 4.10. For the disability rates (Figure 4.9), while the estimates of group $16$ are similar across all models considered, for the other two groups ($23$ and $68$), the three shrinkage fits outperform with similar results the standard and fixed effect fits. The superiority of the shrinkage estimates becomes more evident in the recovery rates (Figure 4.10), where they result in closer approximations to the true rates. This is particularly clear for group $23$, where the fixed effect model estimates a group effect of zero. We also observe the downside of the standard model, which yields identical estimates for all groups. As an additional visual aid to understand the results across all groups, Figure 4.6 presents scatterplots of the (unscaled) estimated group effects under the different models, alongside the corresponding scatterplot of the (true) simulated ones for reference. For the simple model, we observe that several points align along a horizontal line. These correspond to groups without observed disabilities, resulting in absolutely no data on these group’s particular recovery rate. In such a case, ‘full shrinkage’ occurs: the estimated group effect essentially coincides with the mean of the random effect. This phenomenon appears not only for the simple model, but also for the hierarchical and phase-type model; for instance, for the former, the estimated value of $\nu$ is $40.525$, indicating a very small correlation. While the scatterplots are not directly comparable to the simulated values due to scaling, they provide a useful visual representation of the type of dependence each model is able to generate.
Figure 4.6. Scatterplots of simulated and estimated group effects for Case A.
@@
Moving on to Case B, estimates of the disability rates are depicted in Figure 4.9, and a similar conclusion can be drawn to that of Case A, with the shrinkage estimates showing a clear advantage over the standard and fixed effect fits. However, the primary distinction among the shrinkage estimates is observable in the recovery rates, presented in Figure 4.10. More specifically, for groups $16$ and $68$, the models capable of capturing positive correlation – namely, the hierarchical and phase-type models – provide closer approximations to the true disability rates. This leads us to conclude that by allowing dependency between latent group effects, the estimates of recovery and disability rates can mutually benefit. This is confirmed by Figure 4.7, which shows scatterplots of the simulated and estimated group effects. We observe that the hierarchical and phase-type models now produce estimates for groups without observed disabilities aligned along a positively sloped line, as they are able to learn the positive dependence structure. This is in stark contrast to the simple model, which remains unable to do so, resulting in points along a horizontal line.
Figure 4.7. Scatterplots of simulated and estimated group effects for Case B.
Lastly, for Case C, we limit ourselves to presenting the scatterplots of the simulated and estimated group effects, Figure 4.8, where we observe that the phase-type model is able to recover the negative dependence of the group effects. This is something the simple model cannot achieve. Jointly with the results from Cases A and B, this showcases the flexibility of the phase-type model to capture a wide range of dependence structures.
Figure 4.8. Scatterplots of simulated and estimated group effects for Case C.
@@
Figure 4.9. Estimates of disability rates for the three selected groups in Case A and Case B.
@@
Figure 4.10. Estimates of recovery rates for the three selected groups in Case A and Case B.
@@
We conclude this part of the analysis by presenting the loglikelihoods of the fitted mixed Poisson regression models. These loglikelihoods serve as statistical means to further assess the quality of the fits and can be found in Table 4.3. The observations from the table are consistent with what one could expect. In Case A, the simple model outperforms others, with the phase-type model following closely. This result aligns with the nature of the models and the data considered for this case. For Case B, the hierarchical model yields the highest loglikelihood, followed closely by the phase-type model. This outcome is somewhat anticipated since these two models are capable of capturing the positive correlation present in the mixing vector, aligning well with the underlying structure. Finally, for Case C, the phase-type model takes the lead, surpassing the simple model. This is hardly surprising, as it is the only model that can capture a negative correlation. Overall, the phase-type model proves to be very competitive across all cases. Its flexibility to capture different dependence structures highlight its potential as a robust tool for experience rating modeling.
@@
Simple 3924.66-3924.66 3816.12-3816.12 3639.01-3639.01
Hierarchical 3926.73-3926.73 3803.57-3803.57
Phase-type 3925.29-3925.29 3805.56-3805.56 3627.30-3627.30